OpenStudy (anonymous):

Find the seventh term of the following sequence? 1, 2, 4, ...

OpenStudy (anonymous):

2 - 1 = 1 4 - 2 = 2 i am guessing it goes by that pattern so the next number would be x - 4 = 3

OpenStudy (anonymous):

and so on

OpenStudy (callisto):

Another possible one is the sequence of \(2^{n-1}\) ...

OpenStudy (anonymous):

A.12 B.22 C.64

OpenStudy (callisto):

B and C are both possible answers to the question....